# Varieties isogenous to a higher product with prescribed numerical invariants

### v1

@@ Abstract.
The first author acknowledges support from the Deutsche Forschungsgemeinschaft (DFG), grant DZ 105/02. The second author acknowledges support from the Folke Lannér’s Fund and she thanks the Christian-Albrecht University of Kiel for its hospitality.
@@ 1. Introduction
Acknowledgements. We thank Barbara Langfeld for her interest in the subject and her help by confirming several of our results using a computer.
@@ 2. Invariants of varieties isogenous to higher products
For a compact Riemann surface $C$ of genus $g$ the Euler number equals $e(C)=2-2g=-2(g-1)$. As the Euler number is multiplicative with respect to taking products we have $e(C_{1}\times\cdots\times C_{n})=(-2)^{n}\prod_{i=1}^{n}(g(C_{i})-1)$. Finally, since $\pi:C_{1}\times\cdots\times C_{n}\rightarrow X$ is a finite unramified covering, we have $e(C_{1}\times\cdots\times C_{n})=\deg(\pi)e(X)=|G|e(X)$. ∎
@@
Here, the sum runs over all $n$-tuples $(k_{1},\ldots,k_{n})$ of non-negative integers whose sum equals to $1$, i.e. exactly one of the entries equals $1$ and all the others equal 00. Since $H^{0}(C_{i},\mathbb{C})\cong\mathbb{C}$ for all $i$ we obtain
@@
3. Possible ramification structures of genus 00 on finite groups
@@
    $r(i)=3$ if and only if $A_{i}$ belongs to one of the following $23$ triples, with the corresponding $\alpha(A_{i})$ value:
@@ 4. Dimension bounds
From Corollary 3.5 we know that not all $C_{i}$ have genus $g(C_{i})=2$. Now we assume that one of the $C_{i}$, say $C_{1}$, is of genus $2$. As $G$ is absolutely faithful it embeds into $\text{Aut}(C_{1})$. The automorphism groups of surfaces of genus $2$ are all classified by the work of Broughton [3]. From Lemma 4.1 and the classification result [3, Table 4] we then know that $G$ is one of the following groups:
@@
Case (d): $|G|=8$; i.e. $G=Q_{8}$ or $G=D_{4}$. The only vectors of genera to be considered are $(2,\ldots,2,3,5)$ or $(2,\ldots,2,3,3,3)$. As $Q_{8}$ has only one element of order $2$, which is central, condition (3.1) does not hold, as in every possible ramification type there is an element of even order. So we now set $G=D_{4}$. For genus 3, we have $\alpha(A)=4$, and so $A=[2,2,2,2,2]$ or $[2,2,4,4]$. Similarly, for genus 2, we have $A=[2,2,2,4]$ (as no product of 3 elements of order 4 can be trivial), and for genus 5, we obtain $A=[2,2,2,2,2,2]$. Note also that for $[2,2,2,2,2]$, at least one generator is central, as otherwise the total $y$-exponent in any product of the 5 generators is odd. Since the square of an element of order 4 is central, we see that the vector $(2,\ldots,2,3,3,3)$ of genera does not satisfy condition (3.1). So for the remaining case, it suffices to assume that in a generating system of type $[2,2,2,2,2,2]$, none of the generators are central; i.e. every generator is conjugate to $y$ or to $xy$ since there are only two conjugacy classes of non-central elements of order 2 in $D_{4}$. Furthermore, at least one generator is conjugate to $y$ and another generator is conjugate to $xy$, otherwise the system would not generate the group. We claim that condition (3.1) is not satisfied. Indeed, for the generating system of type $[2,2,4,4]$, for the product of the generators, taken in some order to be trivial, the system must be of the corresponding form $[x^{2k+m}y,x^{2\ell+m}y,x^{\epsilon},x^{\delta}]$ for some $k,\ell,m\in\{0,1\}$ and $\epsilon,\delta\in\{\pm 1\}$. Similarly, for genus 2, we have that the corresponding system is of the form $[x^{2k+m}y,x^{2\ell+(m+1)}y,x^{2},x^{\epsilon}]$ for some $k,\ell,m\in\{0,1\}$ and $\epsilon\in\{\pm 1\}$. Finally returning to the system of type $[2,2,2,2,2]$, note that in order to be a generating system at least one generator is conjugate to $y$ and another generator is conjugate to $xy$. Hence the result follows.
@@
$\text{ccl}(x)$ =$=\{x,xy,xw,xw^{2},xyw,xyw^{2}\}$
$\text{ccl}(y)$ =$=\{y\}$
$\text{ccl}(z)$ =$=\{z,yz\}$
$\text{ccl}(w)$ =$=\{w,w^{2}\}$
[…]
$\text{ccl}(yw)$ =$=\{yw,yw^{2}\}$
$\text{ccl}(zw)$ =$=\{zw,yzw^{2}\}$
$\text{ccl}(zw^{2})$ =$=\{zw^{2},yzw\}$
@@ 3. $\bullet$
For each $i\in\{1,\ldots,n\}$ we have $\alpha(A_{i})\leq 84$ (see Corollary 3.4). This implies $|G|^{n-1}\leq 84^{n}$, i.e. $|G|\leq 84^{n/(n-1)}$. On the other hand, as $g(C_{i})>2$ for every $i\in\{1,\ldots,n\}$ from Proposition 4.2, we have $|G|=\prod_{i=1}^{n}(g(C_{i})-1)\geq 2^{n}$. Altogether we obtain $2<84^{1/(n-1)}$. The last expression is decreasing in $n$ and we have $84^{1/7}=1.88\ldots$. This proves that the dimension of $X$ is less or equal to $7$. In the case $n=7$ we must have $g(C_{i})=3$ for all $i\in\{1,\ldots,7\}$. Otherwise we would have $192=2^{6}\cdot 3\leq|G|\leq 84^{7/6}=175.7\ldots$. However there is no group of order $1282^{7}=$ 128 that will satisfy Corollary 3.4. Indeed, if this were the case we would find a spherical system of generators $T$ such that the corresponding ramification type $A$ satisfies $\alpha(A)=2^{6}=64$. But there is no such type in the table of Corollary 3.4. ∎
@@ 5. Possible orders for $G$
For the solvable groups $G$, recall that either the Sylow 2-subgroup is normal or the Sylow 7-subgroup is normal. Indeed, a solvable group has an abelian normal subgroup $N$ of prime order. We may suppose $|N|\neq 7$, else we are done. If $|N|=2$, from considering $G/N$, we deduce that the Sylow 7-subgroup of $G$ is normal. If $|N|=3$, again we are done from considering $G/N$. Now if the Sylow 2-subgroup is normal, we see that no ramification type is possible for an action on a surface of genus 22. So suppose the Sylow 7-subgroup is normal. Consider the quotient of $G$ by its Sylow 7-subgroup. As a spherical system of generators for $G$ is still a spherical system of generators in the quotient, looking at the ramification type $[2,3,7]$, which is the only possibility for an action on a surface of genus 3, if $G$ had a spherical generating system $(g_{1},g_{2},g_{1}g_{2})$ such that $\{\text{o}(g_{1}),\text{o}(g_{2}),\text{o}(g_{1}g_{2})\}=\{2,3,7\}$, then, in the quotient of $G$ by its Sylow 7-subgroup, we would need the generator of order 3 to be the multiplicative inverse of the generator of order 2. Hence, we obtain a contradiction. Thus $G$ here has no action on a surface of genus 3.
@@
$[2,2,2,3],\quad[2,4,12],\quad[2,6,6],\quad[3,6,6],\quad[3,4,4].$
@@
In the first, there are actions on surfaces of genera 3, 4 and 5. In the second, only on surfaces of genera 3 and 4, and in the third, on surfaces of genera 3 and 7. Broughton [3, Table 5] showed that among the groups of order 48, only $Z_{4}^{2}\rtimes Z_{3}$ and $Z_{2}\times S_{4}$ can act on a surface of genus 3, with ramification types $[3,3,4]$ and $[2,4,6]$ respectively.
@@
A direct computation shows that there are only two conjugacy classes of elements of order 3, determined by $x$ and $x^{-1}$ respectively, and each are of size 16. The remaining conjugacy classes consist of one class of elements of order 2 determined by $y^{2}$, and four classes of elements of order 4 determined by $y$, $yz$, $yz^{2}$ and $y^{-1}z$ respectively. Each of these five conjugacy classes is of size 3. Certainly here $G$ also has no element of order 8 or 12. Hence only option (5.1) needs to be considered. For the action on a surface of genus 7, note that the ramification types $[2,2,2,4]$ and $[4,4,4]$ are not possible, since all elements of order 2 and of order 4 lie in the normal subgroup $Z_{4}^{2}$. Hence there is no action on a surface of genus 7. ∎

### v2

@@ Abstract.
The first author acknowledges support from the Deutsche Forschungsgemeinschaft (DFG), grant DZ 105/02. The second author acknowledges support from the Folke Lannér’s Fund and she thanks the Christian-Albrecht University of Kiel for its hospitality. This research was also supported by the Royal Physiographic Society of Lund.
@@ 1. Introduction
Acknowledgements. We thank Barbara Langfeld for her interest in the subject and her help by confirming several of our results using a computer. We also thank the referee for pointing out the work of Polizzi [14], which closes a gap in a result of Broughton [3, Table 5].
@@ 2. Invariants of varieties isogenous to higher products
For a compact Riemann surface $C$ of genus $g$ the Euler number equals e(C)=22g=2(g1)e(C)=2-2g=-2(g-1). As the Euler number is multiplicative with respect to taking products we have $e(C_{1}\times\cdots\times C_{n})=(-2)^{n}\prod_{i=1}^{n}(g(C_{i})-1)$. Finally, since $\pi:C_{1}\times\cdots\times C_{n}\rightarrow X$ is a finite unramified covering, we have $e(C_{1}\times\cdots\times C_{n})=\deg(\pi)e(X)=|G|e(X)$. ∎
@@
Here, the sum runs over all $n$-tuples $(k_{1},\ldots,k_{n})$ of non-negative integers whose sum equals to $1$, i.e. exactly one of the entries equals $1$ and all the others equal 0. Since $H^{0}(C_{i},\mathbb{C})\cong\mathbb{C}$ for all $i$ we obtain
@@
3. Possible ramification structures of genus 0 on finite groups
@@
    $r(i)=3$ if and only if $A_{i}$ belongs to one of the following $22$ triples, with the corresponding $\alpha(A_{i})$ value:
@@ 4. Dimension bounds
From Corollary 3.5 we know that not all $C_{i}$ have genus $g(C_{i})=2$. Now we assume that one of the $C_{i}$, say $C_{1}$, is of genus 22. As $G$ is absolutely faithful it embeds into $\text{Aut}(C_{1})$. The automorphism groups of surfaces of genus $2$ are all classified by the work of Broughton [3]. From Lemma 4.1 and the classification result [3, Table 4] we then know that $G$ is one of the following groups:
@@
Case (d): $|G|=8$; i.e. $G=Q_{8}$ or $G=D_{4}$. The only vectors of genera to be considered are $(2,\ldots,2,3,5)$ or $(2,\ldots,2,3,3,3)$. As $Q_{8}$ has only one element of order $2$, which is central, condition (3.1) does not hold, as in every possible ramification type there is an element of even order. So we now set $G=D_{4}$. For genus 3, we have $\alpha(A)=4$, and so $A=[2,2,2,2,2]$ or $[2,2,4,4]$. Similarly, for genus 2, we have $A=[2,2,2,4]$ (as no product of 3 elements of order 4 can be trivial), and for genus 5, we obtain $A=[2,2,2,2,2,2]$. Note also that for $[2,2,2,2,2]$, at least one generator is central, as otherwise the total $y$-exponent in any product of the 5 generators is odd. Since the square of an element of order 4 is central, we see that the vector $(2,\ldots,2,3,3,3)$ of genera does not satisfy condition (3.1). So for the remaining case, it suffices to assume that in a generating system of type $[2,2,2,2,2,2]$, none of the generators are central; i.e. every generator is conjugate to $y$ or to $xy$ since there are only two conjugacy classes of non-central elements of order 2 in $D_{4}$. Furthermore, at least one generator is conjugate to $y$ and another generator is conjugate to $xy$, otherwise the system would not generate the group. We claim that condition (3.1) is not satisfied. Indeed, for the generating system of type $[2,2,4,4]$, for the product of the generators, taken in some order to be trivial, the system must be of the corresponding form $x^{2k+m}y,x^{2\ell+m}y,x^{\epsilon},x^{\delta}]$ for some $k,\ell,m\in\{0,1\}$ and $\epsilon,\delta\in\{\pm 1\}$. Similarly, for genus 2, we have that the corresponding system is of the form $x^{2k+m}y,x^{2\ell+(m+1)}y,x^{2},x^{\epsilon}]$ for some $k,\ell,m\in\{0,1\}$ and $\epsilon\in\{\pm 1\}$. Finally returning to the system of type $[2,2,2,2,2]$, note that in order to be a generating system at least one generator is conjugate to $y$ and another generator is conjugate to $xy$. Hence the result follows.
@@
$\text{ccl}(x)$ $=\{x,xy,xw,xw^{2},xyw,xyw^{2}\}$
$\text{ccl}(y)$ $=\{y\}$
$\text{ccl}(z)$ $=\{z,yz\}$
$\text{ccl}(w)$ $=\{w,w^{2}\}$
$\text{ccl}(xz)$ $=\{xz,zx,xzw,xzw^{2},zxw,zxw^{2}\}\qquad\text{(order 4)}$
$\text{ccl}(yw)$ $=\{yw,yw^{2}\}$
$\text{ccl}(zw)$ $=\{zw,yzw^{2}\}$
$\text{ccl}(zw^{2})$ $=\{zw^{2},yzw\}$
@@ 3. $\bullet$
For each $i\in\{1,\ldots,n\}$ we have $\alpha(A_{i})\leq 84$ (see Corollary 3.4). This implies $|G|^{n-1}\leq 84^{n}$, i.e. $|G|\leq 84^{n/(n-1)}$. On the other hand, as $g(C_{i})>2$ for every $i\in\{1,\ldots,n\}$ from Proposition 4.2, we have $|G|=\prod_{i=1}^{n}(g(C_{i})-1)\geq 2^{n}$. Altogether we obtain $2<84^{1/(n-1)}$. The last expression is decreasing in $n$ and we have $84^{1/7}=1.88\ldots$. This proves that the dimension of $X$ is less or equal to $7$. In the case $n=7$ we must have $g(C_{i})=3$ for all $i\in\{1,\ldots,7\}$. Otherwise we would have $192=2^{6}\cdot 3\leq|G|\leq 84^{7/6}=175.7\ldots$. However there is no group of order $2^{7}=128$ that will satisfy Corollary 3.4. Indeed, if this were the case we would find a spherical system of generators $T$ such that the corresponding ramification type $A$ satisfies $\alpha(A)=2^{6}=64$. But there is no such type in the table of Corollary 3.4. ∎
@@ 5. Possible orders for $G$
For the solvable groups $G$, recall that either the Sylow 3-subgroup is normal or the Sylow 7-subgroup is normal. Indeed, a solvable group has an abelian normal subgroup $N$ of prime order. We may suppose $|N|\neq 3,7$, else we are done. So $|N|=2$, and from considering $G/N$, we deduce that the Sylow 7-subgroup of $G$ is normal. Suppose first that the Sylow 7-subgroup is normal. Consider the quotient of $G$ by its Sylow 7-subgroup. As a spherical system of generators for $G$ is still a spherical system of generators in the quotient, looking at the ramification type $[2,3,7]$, which is the only possibility for an action on a surface of genus 3, if $G$ had a spherical generating system $(g_{1},g_{2},g_{1}g_{2})$ such that $\{\text{o}(g_{1}),\text{o}(g_{2}),\text{o}(g_{1}g_{2})\}=\{2,3,7\}$, then, in the quotient of $G$ by its Sylow 7-subgroup, we would need the generator of order 3 to be the multiplicative inverse of the generator of order 2. Similarly, if the Sylow 3-subgroup were normal, in the corresponding quotient, we would obtain the generator of order 2 being the multiplicative inverse of the generator of order $7$. Hence, we obtain a contradiction. Thus $G$ here has no action on a surface of genus 3.
@@
$2,2,2,3],\quad[2,4,12],\quad[2,6,6],\quad[3,6,6],\quad[3,4,4].$
@@
In the first, there are actions on surfaces of genera 3, 4 and 5. In the second, only on surfaces of genera 3 and 4, and in the third, on surfaces of genera 3 and 7. From [14, Table 3], which accounts for a missing case in [3, Table 5], among the groups of order 48, only $Z_{4}^{2}\rtimes Z_{3}$, $Z_{2}\times S_{4}$ and the group
$H$ $:=\langle x,y,z,w,t\mid x^{2}=z^{2}=w^{2}=t,\,y^{3}=t^{2}=[x,y]=[x,z]=1,$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad yzy^{-1}=w,\,ywy^{-1}=zw,\,zwz^{-1}=wt\rangle$
can act on a surface of genus 3, with corresponding ramification type $[3,3,4]$, $[2,4,6]$ and $[2,3,12]$ respectively.
@@
A direct computation shows that there are only two conjugacy classes of elements of order 3, determined by $x$ and $x^{-1}$ respectively, and each are of size 16. The remaining conjugacy classes consist of one class of elements of order 2 determined by $y^{2}$, and four classes of elements of order 4 determined by $y$, yzyz, $yz^{2}$ and $y^{-1}z$ respectively. Each of these five conjugacy classes is of size 3. Certainly here $G$ also has no element of order 8 or 12. Hence only option (5.1) needs to be considered. For the action on a surface of genus 7, note that the ramification types $[2,2,2,4]$ and $[4,4,4]$ are not possible, since all elements of order 2 and of order 4 lie in the normal subgroup $Z_{4}^{2}$. Hence there is no action on a surface of genus 7.
So it remains to deal with the case $G=H$. First observe that the following relations follow from the presentation of $G$:
$[x,t]=[z,t]=[w,t]=1.$
We further obtain that $[y,t]=1$. Indeed, from $yz^{3}y^{-1}=w^{3}$ we obtain $yzty^{-1}=wt$. Substituting $yz=wy$ into the previous equation gives $wyty^{-1}=wt$, which then yields $yt=ty$. Lastly, we deduce from the presentation of $G$ that $[x,w]=1$; indeed, from $yz=wy$ we have
$(xwy)^{2}$ $=(xyz)^{2}$
$=x^{2}(yz)^{2}$
$=x^{2}yzy^{-1}y^{-1}z$
$=x^{2}wy^{-1}z.$
As $(xwy)^{2}=xwyxwy=xwxywy$, we obtain from the above that
$wxywy=xwy^{-1}z.$
Recalling that $yz=wy$, we then get
$wxy^{-1}z=xwy^{-1}z,$
and so $wx=xw$, as claimed.
Note also that
$zwz^{-1}=w^{-1},$
and since $z^{2}=w^{2}=t\in Z(G)$, we further have that
$zwz^{-1}=z^{-1}wz,\qquad wzw^{-1}=w^{-1}zw,\qquad zyz^{-1}=z^{-1}yz,\qquad\textup{and}\qquad wyw^{-1}=w^{-1}yw.$
Next, using the relations $yz=wy$ and $yw=zwy$, we compute that
$z^{-1}yz=z^{-1}wy=z^{2}yw=yw^{-1},$
and from $zw=w^{-1}z$,
$w^{-1}zw=z^{-1}.$
Likewise, from $yw=zwy$, we obtain
$w^{-1}yw=w^{-1}zwy=z^{-1}y.$
For convenience, we list the non-trivial conjugacy classes in $G$:
$\textup{ccl}(t)$ $=\{t\}$
$\textup{ccl}(x)$ $=\{x\}$
$\textup{ccl}(xt)$ $=\{xt\}$
$\textup{ccl}(z)$ $=\{z,w,zw,wt,zt,zwt\}\text{ with elements of order }4$
$\textup{ccl}(xz)$ $=\{xz,xw,xzw,xwt,xzt,xzwt\}\text{ with elements of order }2$
$\textup{ccl}(y)$ $=\{y,ywt,zyt=yzwt,wyt=yzt\}\text{ with elements of order }3$
$\textup{ccl}(y^{-1})$ $=\{y^{-1},y^{-1}zw,y^{-1}z,y^{-1}w\}\text{ with elements of order }3$
$\textup{ccl}(yt)$ $=\{yt,yw,yzw,yz\}\text{ with elements of order }6$
$\textup{ccl}(y^{-1}t)$ $=\{y^{-1}t,y^{-1}zwt,y^{-1}zt,y^{-1}wt\}\text{ with elements of order }6$
$\textup{ccl}(xy)$ $=\{xy,xywt,xyzwt,xyzt\}\text{ with elements of order }12$
$\textup{ccl}(xyt)$ $=\{xyt,xyw,xyzw,xyz\}\text{ with elements of order }12$
$\textup{ccl}(xy^{-1})$ $=\{xy^{-1},xy^{-1}zw,xy^{-1}z,xy^{-1}w\}\text{ with elements of order }12$
$\textup{ccl}(xy^{-1}t)$ $=\{xy^{-1}t,xy^{-1}zwt,xy^{-1}zt,xy^{-1}wt\}\text{ with elements of order }12$
As there is no element of order 8 in $G$, there is no action on a surface of genus 4. Indeed, the only possible ramification type for genus 4 is $[2,4,8]$. Therefore we are left with only the option (5.1), with the possible ramification types for a surface of genus 7 being $[2,2,2,4]$ or $[4,4,4]$. Clearly the type $[4,4,4]$ is impossible, as the only elements of order 4 all lie in the union of the conjugacy class of $z$ with $\{x,xt\}$, and $y\notin\langle\textup{ccl}(z)\cup\{x,xt\}\rangle$. Similarly, from the above data, we see that a system of ramification type $[2,2,2,4]$ can never be a generating set for $G$. The proof is now complete. ∎